Proof by contradiction that if n is a positive integer, then n is odd if and only 

if 5n + 6 is odd


1
Expert's answer
2021-04-26T05:59:12-0400

Let us prove by contradiction that if "n" is a positive integer, then "n" is odd if and only if "5n + 6" is odd.


Let "n" is not odd. Then "n" is even, and hence "n=2k,\ k\in\mathbb N." It follows that "5n+6=5(2k)+6=2(5k+3)", and hence "5n+6" is even, that is "5n + 6" is not odd.


On the other hand, let "5n+6" is not odd. Then "5n+6" is even, and hence "5n+6=2t,\ t\in\mathbb N."

It follows that "5n=2t-6=2(t-3)", and hence "5n" is even. Therefore, "n" is also even, that is "n" is not odd.